If 3t - 7 = 5t , then 6t =
Answer by ikleyn(52776)

.
If 3t - 7 = 5t , then 6t = ?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

This problem is solved in two steps.

Step 1.  If  3t - 7 = 5t,  then  -7 = 5t - 3t  and  2t = -7.


Step 2.  Since 2t = -7,  6t is three times (-7), i.e.  6t = 3*(-7) = -21.

Solved.